OpenStudy (anonymous):

FInding the Lambda of a 3x3 matrice.

OpenStudy (anonymous):

\[\left[\begin{matrix} \lambda-4 & 0 & 0 \\ 0 & \lambda & 2 \\ 0 & 3 & \lambda-1\end{matrix}\right]\] find \(\lambda\) !!

OpenStudy (zarkon):

that is a 3x3 matrix. If you are trying to find the eigenvalues then take the determinant and set it equal to zero and solve for \(\lambda\)

OpenStudy (anonymous):

ahh my bad, i typed wrong

OpenStudy (anonymous):

i got \[\lambda^3 - 5\lambda^2-2\lambda+24 = 0\]

OpenStudy (anonymous):

\[\lambda(\lambda^2-5\lambda-2) + 24=0\] what should i do after this?

OpenStudy (anonymous):

find what values for lambda satisfy the equation

OpenStudy (anonymous):

i am stucked

OpenStudy (anonymous):

so lambda = ?, ??, ??? A third order polynomial should yield 3 roots, though they might not be distinct. Those are what your lambda values are

OpenStudy (anonymous):

x(x^2−5x−2)+24=0 Find the roots

OpenStudy (anonymous):

wait a minute

OpenStudy (anonymous):

λ = -2,3,4 Those are your eigenvalues

OpenStudy (anonymous):

\[\lambda (\lambda^2 - 5\lambda - 2 + 8 - 8) + 24 = 0 \] \[\lambda(\lambda^2 - 5\lambda + 6 - 8) + 24 = 0\] \[\lambda([(\lambda -2) (\lambda - 3)] - 8) + 24 = 0\]

OpenStudy (anonymous):

\[\lambda((\lambda-2)(\lambda-3) - 8) = -24\]

OpenStudy (anonymous):

You have to multiply everything out, and combine terms of the same power, then factor

OpenStudy (anonymous):

i took a shortcut http://www.wolframalpha.com/input/?i=x%28x^2%E2%88%925x%E2%88%922%29%2B24%3D0

OpenStudy (anonymous):

ahh..,

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

thank you !!